OpenStudy (anonymous):

solve the following system by using row-equivalent matrices x+4y+z=18, 3x+3y-2z=2, -4y+z=-7

OpenStudy (unklerhaukus):

\[\left(\begin{array} \ 1&4&1\\3&3&-2 \\0&-4&1\end{array}\right) \left(\begin{array} \\x\\y\\z\\ \end{array}\right)=\left(\begin{array} 18\\2\\-7\\ \end{array}\right)\]

OpenStudy (anonymous):

use cramer

OpenStudy (anonymous):

well i don't think it is the row- equivalent matrices

OpenStudy (anonymous):

well i did use the cramer's rule and i also got the answer but it is needed to be done using row-equivalent matrices

OpenStudy (anonymous):